y varies inversely with x and y = 8 when x = 2.

Which is the inverse variation equation for the relationship?




A.

Y=16/X


B.

y = 2x


C.

y = 16x


D.

y = x + 2

Answer by ewatrrr(24785)

 
Hi,
y = k/x
8 = k/2
16= k
y = 16/x